# Co-Designing Multimodal Systems for Accessible Remote Dance Instruction

### v1

@@ 4.2. Modalities for Dance Instruction
Participants used a diversity of strategies to convey motion in the small group sessions (teaching phase, designing phase): verbal descriptions, non-verbal audio, and haptic or tactile approaches.
@@ 4.2.1. Verbal Descriptions
G5, G6, G8 provided high-level, cross-movement context, either before or during instruction of the routine. For example, before teaching, A8 explained the cha-cha clip by saying, “The dance is working in opposites […] so if you go forward, you will complement that by going backwards.” B8a noted that descriptions should first explain the rhythm using verbal counts before moving into defining the steps. On the other hand, B5a and B5b described the rhythm after movement definitions. D6 noted how it was better to use semantic cues that reference movements that happen at a specific rhythm, such as ‘forward’, ‘back’, and ‘close’, instead of counts.
@@
Takeaway: Provide verbal feedback by offering incremental corrections and affirmations. Adapt verbal descriptions based on the learner’s correctness and understanding of current descriptions.
@@
Learner familiarity with the dance routine or style determines whether to use full descriptions or movement vocabulary terms as a shorthand. G2, G4, G5, G6, G7, and G8 varied the descriptiveness of verbal instructions, either as the learner got more advanced or based on their level of experience. For example, G4 noted that it would be efficient to convert to shorter phrases once learners have understood the motions, such as: ’arms to side’ and ’bend knees’. On the other hand, B8a was already familiar with a ‘Triple Step’ in cha-cha and did not need that movement defined.
G5 and G8 found it helpful to repeat words and phrases as students learned and practiced movements. B6, for example, expressed how repetition helps students remember steps and correct mistakes: “I personally like [redundancy]. The more you say it, the more it sticks. Keep saying ’left’, keep saying ’right”’
@@ 4.2.3. Haptic and Tactile Input
Takeaway: Haptics for remote learning can function as a parallel to touch feedback to instruct foot positioning, body weight distribution, and movement duration.
@@ 4.3. Multimodal Systems for Dance Instruction
G1 and G8 provided movement vocabularies that matched all movement descriptions to parallel verbal names, sound cues, and haptic cue patterns (Figure 6, Group 1 and Figure 7, Group 5). G2, G3, G4 (Figure 6), G5, G7, and G8 (Figure  7) included verbal names and sound cues in their movement vocabularies.
@@ 4.3.2. Phases of dance instruction
G2, G3, G4, and G5 noted that it is important to have specificity on the structure of the movements when first learning and less emphasis on feelings and textures in these early stages. Later, emotions, textures, and how to move would be introduced with non-verbal audio once dancers have practiced what to do. As this occurs, verbal descriptions also become less detailed. For example, after B8a Rock Steps and Triple Steps in detail, she described the routine as: “Da da, cha cha cha” to establish the rhythm, where da represented rocks and cha represented parts of the Triple Step.
@@
G5 discussed how these phases can differ based on prior knowledge. B5b explained, “I think it depends on your experience, because I’m coming from no experience. So the importance to me is to let me know how to move my body.” However, more experienced dancers may prioritize expressiveness as they are already familiar with movement structure.
@@ 4.3.3. Tradeoffs Between Modalities
Sound cues were viewed as faster and more efficient than verbal descriptions, especially when overlaid with dance music. G1, G2, G4, and G8 described how learners could “tune out words” (P1) but process sound cues automatically once mappings were learned. Participants expressed that sound made the experience feel more natural and engaging, with P4 explaining, “It’s nicer to dance to music.” However, G3, G6, and G7 cautioned that mapping too many motion parameters to sound increased cognitive load. G6 concluded that simple mappings, such as pitch to height, were preferable in instructional contexts, even if they sacrificed representational accuracy.
Haptics were described by G1, G3, G7, and G8 as efficient and immediate, with learners able to process cues such as left–right vibrations more quickly than directional words. Participants emphasized that haptic cues also build muscle memory, allowing movements to become more internalized over time (P8a). In partner dance contexts, D8b noted that tactile input is essential for coordination and connection, which G8 sought to replicate through haptic design. Still, groups identified limits: G2 considered haptics overwhelming for contemporary dance with too many degrees of freedom, while G4 described haptic guidance as trial-and-error for complex motions. G8 also stressed cognitive limits, noting that learners could process only a few simultaneous taps before input became confusing, motivating them to restrict the number of haptic devices in their design.
@@ 4.3.5. Customization
G4 noted that differences in skill levels of learners may lead to differences in ’correctness’ criteria for feedback systems. For instance, since different users may have differences in flexibility, they determined that the system may or may not check if the dancer’s hands touched the ground during the lunge step when providing feedback.
G5’s system (Figure 7, Group 5) was described as a ‘Choose Your Own Adventure’ game where users had different settings options, including level of detail in verbal descriptions, speed of the dance, adding a metronome beat, when to include sound cues, and when to include music. For beginners, G5 concluded that it is better to have verbal descriptions only, and gradually add non-verbal elements as the user becomes more proficient. However, for advanced learners, they preferred to have non-verbal elements incorporated at the beginning to establish fluidity, expressiveness, and tempo. G5 additionally noted that people have different preferences between auditory and tactile input.
@@ 5. Discussion
We synthesize design guidelines for potential future systems, reflect on limitations of our workshop, and propose directions for future work.
5.1. Design Guidelines
We summarize major results and share 20 design guidelines for future remote multimodal dance instruction systems (Table 5).
Guideline Description
DG1 Create a multimodal movement vocabulary to break down pieces dance routines using memorable movement names or numbers to reference movements.
@@ 5.1. Design Guidelines
DG2 Break individual movement instruction into specific body part positioning, overall body positioning and structure, and body parts in relation to one another.
DG3 Use detailed verbal descriptions for movement structure, orientation, and weight distributions.
@@
DG4 Provide high-level, cross-movement context on dance routine patterns and rhythm.
DG5 Use learning phases to break down descriptions and let users get accustomed to cue mappings.
DG6 Allow navigable steps that users can repeat for practice.
DG7 Incorporate practice phases that emphasize timing, add dance music, and reduce verbal cues.
DG8 Use haptic and tactile elements to provide immediate spatial and orientation feedback.
DG9 Use sound cues for positive affirmations to provide encouragement.
DG10 Allow verbal check-ins and incremental corrections to support dialogue and learner agency.
DG11 In dances where limb positions matter, use multimodal sound and haptic techniques to convey learners’ live limb positions.
DG12 Allow customizable experiences that account for prior knowledge of dance style and skill level.
DG13 Enable learners to configure systems based on modality preferences (auditory vs. tactile).
@@
DG14 Use semantic sound-to-movement mappings to convey fluidity, weight, and expressiveness.
DG15 Use direct sound-to-movement mappings to convey movement structure and duration.
DG16 Use tactile metaphors and other music snippets to convey expressive movement qualities.
DG17 Use non-verbal sound cues and verbal counts to establish tempo.
DG18 Haptic and tactile feedback is important for systems teaching partner dance styles.
DG19 Haptic feedback is useful for conveying repetitive and rhythmic motions.
DG20 Sound cues and tactile metaphors are important for teaching expressive dance forms.
Table 5. Design guidelines for accessible remote dance instruction systems. Design guidelines related to breaking down movements (DG1-DG4), using instruction phases to build familiarity (DG5-DG7), providing feedback with multiple modalities (DG8-DG11), allowing customization (DG12, DG13), using multiple modalities and language techniques to convey expressive movement (DG14-DG17), and considering unique characteristics of dance styles (DG18-DG20).
5.1.1. Breaking Down Instruction.
@@
We observed that groups provided high-level context for routines, broke routines down into movements, and further deconstructed those movements into lower-level movement descriptions (DG1-DG4). A common existing approach to dance instruction for all learners is breaking down routines into more manageable pieces, such as into eight-counts. Our workshops revealed that BLV students required more fine-grained dissecting of dance routines into distinct portions for learning movement structure. This is accomplished through multi-modal movement vocabularies, and further breaking down of individual movements with local and overall body descriptions and the relationship between body parts. As BLV learners can not directly preview the video, providing high-level context to frame the dance is also valuable.
@@
In our study, participants typically used their own vocabularies to name and describe movements, occassionally using an existing movement term (e.g., triple-step). However, prior work has proposed structured frameworks to describe movements. For example, the Locus framework  (Graham, 2016; Sulzman, 1978) maps reference points in a body-sized cube to numbers or letters of the alphabet that are consistent across performances. To reduce cognitive load, our participants described motions locally—such as starting with the legs, then progressing to the torso, arms, and head—and separately combining them into a full movement (DG2). Similarly, Laban Movement Analysis (LMA)  (Groff, 1995) explicitly categorizes movements based on what parts are moving, orientation and direction of these parts, overall body shape, and dynamics such as weight and flow. To increase consistency between dances, instructors or future systems could use such frameworks for movement instruction.
@@
Further, BLV participants raised common stategies from their lived experience for conveying orientation and spatial information such as using the clock to convey orientation (Tabb, 2025). Thus, future instructors or systems could tailor their descriptions using best practices from orientation and mobility training or BLV learners’ own preferred terminology.
5.1.2. Instructional Phases.
@@
Accessible remote dance instruction systems should teach in phases (DG5-DG7). Our results conclude that early instructional phases should prioritize detail in vocabulary definitions, verbal descriptions, and specific feedback as users internalize movement structure. Following this, users are able to build practice through repetition and the emphasis on non-verbal cues to establish timing and expressive qualities. Eventually, users are able to explore how the routine interacts with music by overlaying the music on non-verbal cues, before eventually removing all cues. Repetition was identified as essential for BLV learners to develop body awareness and internalize choreography, making structured practice phases (DG7) crucial for dance instruction systems. Our results found that repeating individual steps while learning a full routine is important for building confidence. Systems should allow for navigable steps that users can revisit and practice while putting routines together (DG6). These learning phases reflect the Fitts and Posner model for motor skill learning, where learners gradually increase autonomy in motor skills with practice (Fitts and Posner, 1967). Insights on modality usage taken from our work can be used to extend this model into accessibility contexts to provide BLV learners with greater confidence and the practice required for autonomy in movement.
5.1.3. Feedback
@@
Our results showed that BLV learners benefit from immediate feedback to confirm orientation, spatial positioning, and overall confidence in their movements. Thus, we derived multimodal feedback techniques (DG8-DG11). Our findings resonate with Hattie and Timperley’s Feedback Model (Hattie and Timperley, 2007) for learning which outlines three questions that effective feedback should ask: Where am I going? (goals for movement), How am I going? (current progress towards goals), and Where to next? (corrections needed to reach goals). Future accessible dance instruction systems can incorporate this model into its feedback technique by providing clear descriptions of movement goals (DG3), multimodal affirmations on movement correctness (DG9, DG11), and incremental feedback to correct positioning and orientation (DG8, DG10).
5.1.4. Customization
@@
As outlined by participants and prior work (Aflatoony and Lee, 2020; Jones et al., 2024), customizable technologies improve the technology experience of people with disabilities. Our results showed clear differences across learners: some valued detailed auditory explanations, while others preferred tactile or haptic cues that allowed them to physically establish orientation and weight. Users additionally have varied requirements in the level of detail required in descriptions. Thus, remote education systems should allow customizable experiences that acccount for prior knowledge of dance style and skill level (DG12) and enable learners to configure systems based on modality preferences (auditory vs. tactile) (DG13).
5.1.5. Temporal and Expressive Qualities
Many dance styles contain temporal (speed-based) and expressive (feeling-based) qualities that are critical for dance instruction (Blom and Chaplin, 2004). Non-verbal sound cues are an efficient way to convey such aspects that are difficult to verbalize concisely. Cues can be mapped semantically (DG14) or directly (DG15) to motions. Tactile metaphors and music clips can convey expressive movement qualities (DG16), while non-verbal sound cues can be used to establish tempo (DG17). Our results, aligning with prior work (De Silva et al., 2023; Bläsing and Zimmermann, 2021), highlights the importance of sound cues and metaphors to convey nuanced movement qualities to BLV individuals. LMA’s Effort category captures temporal and expressive qualities of movements, such as Weight (strong vs. light), Flow (bound vs. free), Time (sudden vs. sustained), and Space (direct vs. indirect). This provides a systematic vocabulary for such dimensions that may be useful for future systems.
5.1.6. Differences Between Dance Styles

### v3

@@ 4.2. Modalities for Dance Instruction
Participants used diverse strategies to convey motion in the small group sessions (Teaching phase, Designing phase): verbal descriptions, non-verbal audio, and haptic or tactile approaches.
@@ 4.2.1. Verbal Descriptions
G5, G6, G8 provided high-level,
cross-movement context, either before or during instruction of the routine. For example, before teaching, A8 explained the cha-cha clip by saying, “The dance is working in opposites […] so if you go forward, you will complement that by going backwards.” B8a noted that descriptions should first explain the rhythm using verbal counts before moving into defining the steps. On the other hand, B5a and B5b described the rhythm after movement definitions. D6 noted how it was better to use semantic cues that reference movements that happen at a specific rhythm, such as ‘forward’, ‘back’, and ‘close’, instead of counts.
@@
Takeaway: Provide verbal feedback by offering incremental corrections. Adapt verbal descriptions based on the learner’s correctness and understanding of current descriptions.
@@
Learner familiarity with the dance routine or style determines whether to use full descriptions or movement vocabulary terms as a shorthand. G2, G4, G5, G6, G7, and G8 varied the descriptiveness of verbal instructions, either as the learner got more advanced or based on their level of experience. For example, G4 noted that it would be efficient to convert to shorter phrases once learners have understood the motions, such as: ‘arms to side’ and ‘bend knees’. On the other hand, B8a was already familiar with a ‘Triple Step’ in cha-cha and did not need that movement defined.
G5 and G8 found it helpful to repeat words and phrases as students learned and practiced movements. B6, for example, expressed how repetition helps students remember steps and correct mistakes: “I personally like [redundancy]. The more you say it, the more it sticks. Keep saying ‘left’, keep saying ‘right”’
@@ 4.2.3. Haptic and Tactile Input
Takeaway: Haptics for asynchronous learning can function as a parallel to touch feedback to instruct foot positioning, body weight distribution, and movement duration.
@@ 4.3. Multimodal Systems for Dance Instruction
G1 and G5 provided movement vocabularies that matched all movement descriptions to parallel verbal names, sound cues, and haptic cue patterns (Figure 7, Group 1 and Figure 7, Group 5). G2, G3, G4 (Figure 7), G5, G7, and G8 (Figure  7) included verbal names and sound cues in their movement vocabularies.
@@ 4.3.2. Phases of dance instruction
G2, G3, G4, and G5 noted that it is important to have specificity on the structure of the movements when first learning and less emphasis on feelings and textures in these early stages. Later, emotions, textures, and how to move would be introduced with non-verbal audio once dancers have practiced what to do. As this occurs, verbal descriptions also become less detailed. For example, after B8a described Rock Steps and Triple Steps in detail, she described the routine as: “Da da, cha cha cha” to establish the rhythm, where da represented rocks and cha represented parts of the Triple Step (Figure 7, Group 8).
@@
G5 discussed how these phases can differ based on prior knowledge (Figure 7, Group 5). B5b explained, “I think it depends on your experience, because I’m coming from no experience. So the importance to me is to let me know how to move my body.” However, more experienced dancers may prioritize expressiveness as they are already familiar with movement structure.
@@ 4.3.3. Tradeoffs Between Modalities
Sound cues were viewed as faster and more efficient than verbal descriptions, especially when overlaid with dance music. G1, G2, G4, and G8 described how learners could “tune out words” (B1) but process sound cues automatically once mappings were learned. Participants expressed that sound made the experience feel more natural and engaging, with B4 explaining, “It’s nicer to dance to music.” However, G3, G6, and G7 cautioned that mapping too many motion parameters to sound increased cognitive load. G6 concluded that simple mappings, such as pitch to height, were preferable in instructional contexts, even if they sacrificed representational accuracy.
Haptics were described by G1, G3, G7, and G8 as efficient and immediate, with learners able to process cues such as left–right vibrations more quickly than directional words. Participants emphasized that haptic cues also build muscle memory, allowing movements to become more internalized over time (B8a). In partner dance contexts, D8b noted that tactile input is essential for coordination and connection, which G8 sought to replicate through haptic design (Figure 7, Group 8). Still, groups identified limits: G2 considered haptics overwhelming for contemporary dance with too many degrees of freedom, while G4 described haptic guidance as trial-and-error for complex motions. G8 also stressed cognitive limits, noting that learners could process only a few simultaneous taps before input became confusing, motivating them to restrict the number of haptic devices in their design.
@@ 4.3.5. Customization
Systems offered areas of customization based on individual needs and dance styles. G4 noted that differences in skill levels of learners may lead to differences in ‘correctness’ criteria for feedback systems. For instance, since different users may have differences in flexibility, they determined that the system may or may not check if the dancer’s hands touched the ground during the lunge step when providing feedback (Figure 7, Group 4).
G5’s system (Figure 7, Group 5) was described as a ‘Choose Your Own Adventure’ game where users had different settings options, including level of detail in verbal descriptions, speed of the dance, adding a metronome beat, when to include sound cues, and when to include music. For beginners, G5 concluded that it is better to have verbal descriptions only, and gradually add non-verbal elements as the user becomes more proficient. However, for advanced learners, they preferred to have non-verbal elements incorporated at the beginning to establish fluidity, expressiveness, and tempo. G5 additionally noted that people have different preferences between auditory and tactile input. Similarly, G8 concluded that less descriptiveness was required for people with more dance experience. Their system design (Figure 7, Group 8) involved a customization step as well, where users can customize level of detail and specific haptic cue patterns.
Teaching techniques and correctness criteria additionally required customization according to dance style. Social dance groups, such as G6 and G8 (Figure 7), emphasized correctness in timing and direction over specific movement mechanics. On the other hand, D6 explained that certain competitive styles emphasized exact movement structure. This led to differences in required cues from the system, such as sound cues conveying exact limb positions:
“In competition ballroom dance, the quality and exactly how you do something matters. In social dance, it’s as long as you and your partner are having fun and you’re getting the steps in time […] So we can have the pitch [of the sound cue] indicate the height [of the hand], but does it matter for this dance?”
D5 additionally explained that for Contemporary dance, applying personalized styles and textures is important. Therefore, it was critical for their system to empower BLV learners to express themselves by conveying more interpretive elements, such as movement fluidity and emotions (Figure 7, Group 5).
@@ 5. Discussion
We synthesize design implications for potential future systems, reflect on limitations of our workshop, and propose directions for future work.
5.1. Design Implications
From participants’ system designs, we created 13 design implications (DI1-DI13) for future asynchronous multimodal dance instruction systems (Table 5). These implications address five system goals: convey movement structure, convey timing and expressiveness, instruct in phases, provide feedback, and provide customization.
System Goal Implication Modalities Description Groups Related Work
Convey movement structure DI1 V, S, H Create a multimodal movement vocabulary to break down dance routines into individual, named movements. 1, 2, 3, 4, 5, 6, 7, 8 (Endo et al., 2024),(Bennett, 2008),(Hassan et al., 2024), (Bläsing, 2015),(Atã and Queiroz, 2019),(Puri, 2004)
@@ 5.1. Design Implications
DI2 V Break individual movements down into verbal descriptions of specific body part motions, body parts in relation to one another, and high-level body positioning and structure. 2, 3, 4, 5, 8 (Rector et al., 2013),(De Silva et al., 2025),(El Raheb et al., 2018)
@@
DI3 V Verbally provide high-level, cross-movement context on dance patterns. 6, 7, 8 (Van Daele et al., 2024)
DI4 H Use haptics to indicate direction of movements 1, 6, 8 (Aggravi et al., 2016),(Hong et al., 2017),(Teng et al., 2025)
@@
Convey timing & expressiveness DI5 V, S Use direct and semantic sound-to-movement mappings and tactile metaphors to convey fluidity, speed, and expressiveness. 1, 2, 3, 4, 5, 7 (De Silva et al., 2025), (Oppici et al., 2020)
DI6 V, S Use non-verbal sound cues and verbal counts to establish tempo. 2, 5, 8 (Keevallik et al., 2024),(Lee et al., 2024)
DI7 H Use haptics to indicate timing of movements. 1, 6, 7, 8 (Aggravi et al., 2016),(Morelli et al., 2010b),(Holland et al., 2018)
Instruct in phases DI8 V, S, H Use learning phases to let users learn movement structure with navigable steps that can be repeated as necessary. 1, 2, 4, 5, 6, 7, 8 (Ulfa and Narawati, 2020),(Hassan et al., 2024), (Fitts and Posner, 1967)
DI9 S, H Use practicing phases to let users practice movements, learn timing, and incorporate dance music. 1, 2, 4, 5, 6, 7, 8 (Zhou et al., 2021)
Provide feedback DI10 H Use haptic and tactile elements to provide spatial and orientation feedback. 1, 3, 4, 6, 7, 8 (Morelli et al., 2010a),(Seham and and, 2015)
DI11 S Use sound cues to provide movement structure correctness feedback. 6, 8 (Oh et al., 2013),(Françoise et al., 2016)
DI12 V Include verbal check-ins and incremental corrections. 2, 4, 5, 6, 8 (Huh et al., 2025),(Ning et al., 2025)
Provide customization DI13 V, S, H Allow customization based on prior knowledge of dance and skill level. 4, 5, 6, 8 (Blanchet et al., 2025)
DI14 V, S, H Customize system based on dance style being taught. 5, 6, 7 (Lee et al., 2024),(Conroy and Bresnahan, 2025)
Table 5. Design implications for accessible remote dance instruction systems categorized into core system goals, including modalities addressed, participant groups that incorporated each guideline, and related work supporting each guideline. V = verbal, S = sound (non-verbal), H = haptic/tactile. Citations marked with are accessibility-related, are dance-related, and are non-accessibility work using the technique in other tasks.
5.1.1. Convey movement structure.
@@
A common existing approach to dance instruction for all learners is breaking down routines into more manageable pieces, such as into eight-counts (Bläsing, 2015). In our study, all groups similarly segmented routines into individual movements (DI1). While sighted learners might rely on visual demonstrations to learn each segmented movement (Bläsing et al., 2018; Bläsing, 2015), BLV learners in our workshops further incorporated a second level of segmentation with verbal descriptions of individual body parts and what they were doing (DI2). As BLV learners can not directly preview the video or instructor, providing high-level context to frame (e.g., “this dance works in opposites; every time you step forward and vice versa”) the dance is also valuable (DI3).
@@
In our study, participants typically used their own vocabularies to name and describe movements, occassionally using an existing movement term (e.g., triple-step, Plié). Dance styles such as Ballet, Cha-cha, and Hip-Hop contain standardized, named movements or sequences (Atã and Queiroz, 2019; Bennett, 2008) that can be used in future systems. Similarly, Laban Movement Analysis (LMA)  (Groff, 1995) explicitly categorizes movements based on what parts are moving, orientation and direction of these parts, overall body shape, and dynamics such as weight and flow. To increase consistency between dances, instructors or future systems may use such existing dance vocabularies or frameworks to name movements.
@@
Furthermore, BLV participants raised common stategies from their lived experience for conveying orientation and spatial information such as using the clock to convey orientation (Tabb, 2025). Thus, systems could tailor their descriptions using best practices from orientation and mobility training or BLV learners’ own preferred terminology.
5.1.2. Convey Timing and Expresssiveness
Many dance styles contain temporal (speed-based) and expressive (feeling-based) qualities that are critical for dance instruction (Blom and Chaplin, 2004). Among our participants, non-verbal sound cues and metaphors proved to be a popular way to convey such aspects that are difficult to objectively describe. Existing movement analysis work (Frid et al., 2018; Landry and Jeon, 2020) shows that sound effectively conveys temporal and expressive movement qualities. Our BLV participants emphasized having sound-to-movement mappings in their dance instruction systems as an efficient way of grasping these movement qualities that sighted learners are able to perceive visually — for example, the duration of the sound cue matching movement duration. Prior work on accessible dance (De Silva et al., 2023; Bläsing and Zimmermann, 2021) additionally highlights the importance of using metaphors to verbally convey these nuanced movement qualities to BLV individuals.
The use of non-verbal sound beats and verbal counts to indicate rhythm (DI6) is also prevalent in existing dance literature (Keevallik et al., 2024; Lee et al., 2024). Our workshops showed that these techniques provide additional grounding for BLV participants who rely more heavily on auditory information in place of visual demonstrations.
5.1.3. Instruct in Phases.
@@ 5.1.2. Convey Timing and Expresssiveness
Accessible remote dance instruction systems should teach in phases (DI8-DI9). Our results conclude that early instructional phases should prioritize detail in vocabulary definitions, verbal descriptions, and specific feedback as users internalize movement structure. Following this, users are able to build practice through repetition and the emphasis on non-verbal cues to establish timing and expressive qualities. Users are then able to explore how the routine interacts with music by overlaying the music on non-verbal cues, before eventually removing all cues. BLV participants emphasized repetition as important for developing body awareness and internalize choreography (DI8). Systems should thus allow for navigable steps that users can revisit and repeat while putting routines together. These instructional phases reflect existing methods for motor skill learning, where learners gradually increase autonomy with repetition and practice (Hassan et al., 2024; Fitts and Posner, 1967; Ulfa and Narawati, 2020). Our work shows that such repetition and practice is particularly important in accessible dance instruction systems to allow BLV learners to form accurate mental models of spatial information that they are not able to visually perceive.
5.1.4. Provide Feedback
@@
Our results showed that BLV learners benefit from immediate feedback to confirm orientation, spatial positioning, and overall confidence in their movements. Thus, we derived multimodal feedback techniques (DI10-DI12). Movement correction and feedback is widely used in existing dance practice. However, the ambiguity of verbal descriptions is higher for BLV learners (De Silva et al., 2025; Bläsing and Zimmermann, 2021), while sighted learners are able to implicitly receive feedback through visual comparison of their movements (Bläsing et al., 2018). BLV participants thus emphasized having clarity in their movement goals and frequent feedback on their movements. Future accessible dance instruction systems can incorporate this model by providing clear descriptions of movement goals (DI2), multimodal feedback on movement correctness (DI10-DI12), and incremental feedback to correct positioning and orientation (DI12).
5.1.5. Provide Customization
@@
As outlined by participants and prior work (Aflatoony and Lee, 2020; Jones et al., 2024; Kane et al., 2014), customizable technologies improve the technology experience of people with disabilities. Our results showed clear differences across learners: some valued detailed auditory explanations, while others preferred tactile or haptic cues that allowed them to physically establish orientation and weight. Users additionally had varied level of detail required in descriptions. For example, participants B5a, B6, B7b and B8 had prior dancing experience, allowing them to better follow the dance to the point where they could teach other participants. Comparatively, BLV participants with no experience required more explicit instructions and greater detail to learn the same dances. Asynchronous dance instruction systems should thus allow for customizable experiences that acccount for prior knowledge of dance style and skill level (DI13). Our workshops additionally revealed the need for system differences across dance styles (DI14), which is discussed in section 5.2.
5.2. Applying Design Implications to Asynchronous Dance Instruction Systems for BLV Learners
To support researchers and practitioners applying the co-designed system architectures and design insights in their work, we share example dance instruction systems, and describe how design implications may vary across dance styles.
5.2.1. Example Accessible Asynchronous Dance Instruction Systems
Our two hypothetical examples of accessible systems each incorporate a subset of our design implications to guide future systems, derived from participant co-designed systems. One example is for a system addressing expressive dance routines, such as with Ballet, Hip-hop, and Contemporary dance. The second example is a virtual dance partner to teach BLV users social dance styles.
Expressive Dance Routine Instruction System: Informed by G4 and G5 designing systems for Contemporary dance routines, this system example teaches a choreographed routine (a set of predefined movements) that incorporates stylistic elements such as weight, fluidity, and expressiveness.
In a preprocessing phase, the system takes a video of a choreographed routine and segments it into a set of individual movements (DI1) (e.g., using a Temporal Convolutional Network to segment distinct dance movements as in prior work (Endo et al., 2024)). Next, the system provides detailed text descriptions of each movement (DI2) (e.g., with a motion-specific vision language model (Chen et al., 2024)). For each segmented movement, the system also creates sound-to-movement mappings (DI5). Direct mappings are created by using pose detection models such as MediaPipe (Sengar et al., 2024; Lugaresi et al., 2019) and AlphaPose (Fang et al., 2022) to extract joint data and sonify them (e.g., mapping pitch to height or volume to speed). To create semantic mappings, the system may first generate expressive descriptions of each movement (Yazdian et al., 2025) and then create expressive sound cues from these descriptions with a text-to-audio model (Huang et al., 2023). During system setup, users may include information on prior dance experience to provide a customized learning experience (DI13) — e.g., the system may substitute detailed low-level descriptions for familiar dance terms for if the participant is familiar with the dance style.
Participants can then use the system by freely navigating via screen reader or voice command between different movements to learn and practice each motion by listening to the verbal and sound cues (DI8). Once a user is comfortable with steps, they may practice them at full speed and on top of existing dance music during practice phases (DI9). Throughout these phases, the system will further collect a video stream of the user’s attempts for each step and provide verbal corrections (DI12) and positive sound cues once a user performs a step correctly (DI11).
Virtual Dance Partner System: Informed by G1, G6 and G8 designing systems for Salsa and Cha-cha, we envision a virtual dance partner for teaching steps from social dance videos that often demonstrate repeating standardized steps. Similar to G8, this system will teach a set of standardized steps (e.g., rock-step, triple-step) (DI1) using verbal descriptions (D12) combined with a pair of haptic probes on the learner’s back that will vary haptic vibration intensity and location corresponding to weight distribution and direction of foot placement (DI4) (e.g., similar to body worn haptics for acrobatics (Popp et al., 2023)). The system will include wrist-worn haptics (Hong et al., 2017) that vibrate to cue movements (DI4, DI7) (i.e., left wrist vibrating means to step left, a double vibration means to turn). Once learning the named movements, users are able to practice these steps with haptic cues (DI9).
5.2.2. Application Across Dance Styles
Dance encompasses a wide range of styles differing in movement standardization and learning techniques (Conroy and Bresnahan, 2025; Foster, 1986; Nicholas and Morris, 2004). Certain styles emphasize standard positions and sequences (e.g., ballet (Bennett, 2008), Indian classical dance (Puri, 2004)). Other styles like Butoh (Kasai, 1999), Gaga (Katan-Schmid, 2016), and Contemporary Improvisation (Nakano and Okada, 2012) prioritize sensation, and internal imagery over the reproduction of fixed shapes (Lepecki, 2006). Instruction for these dance styles rely on sensory descriptions to guide dancers (Seham and and, 2015), offering more inclusive learning by de-emphasizing the reliance on visual replication. In this work, we address asynchrnous learning from online dance tutorial videos, which predominantly feature structured movement routines and timing. This sequence-based, visually taught format presents distinct challenges for BLV audiences. However, we acknowledge the broader landscape of more accessible, improvisational dance practices that are less represented in such online dance tutorials.